OpenStudy (anonymous):

what is the equation of a line with a slope of -3 that passes through the point (3,-3)?

OpenStudy (anonymous):

We can write the equation of the line as: \[(y-b)=m(x-a)\] where m = the slope, and a and b are the coordinates (a,b)

OpenStudy (anonymous):

So all that you have to do is substitute: m = -3 a = 3 b = -3 into: (y-b)=m(x-a)